OpenStudy (anonymous):

The length of a rectangle exceeds its width by 4 meters and its perimeter is 68 meters. Find the dimensions of the rectangle.

OpenStudy (anonymous):

32 and 36

OpenStudy (anonymous):

answer^^^^ link

Parth (parthkohli):

Let us assume that the length is x + 4 and the width is x. \( \color{Black}{\Rightarrow 2(l + b) = P}\) \( \color{Black}{\Rightarrow 2(x + 4 + x ) = P}\) \( \color{Black}{\Rightarrow 2(2x + 4) = P}\)

OpenStudy (anonymous):

let the two consecytive even number be x and x+2 x+x2=150 2x+2=150 2x=148 x=74 the two numbers are 74and 76

OpenStudy (anonymous):

*consecutive

OpenStudy (anonymous):

here is a simple solution: given that there are two quantities involved, divide the perimeter by two (68/2=34) then divide the exceed of the other number to the other number by 2. (4/2=2) then, add and subtract 2 from 34 (34-2=32) (34+2=36) answers: 32 and 36

OpenStudy (anonymous):

the answer is 74 and 76

OpenStudy (anonymous):

that's for the first question. lolz

OpenStudy (anonymous):

:D